Susan, a single taxpayer, owns and operates a bakery as a sole proprietorship. The business is not a specified services business. In 2019, the business pays $60,000 of W–2 wages and reports qualified business income of $200,000. Susan also has a part-time job earning wages of $11,000 and receives $3,200 of interest income. Assume the QBI amount is net of the self-employment tax deduction.

What is Susan's tentative QBI based on the W–2 Wages/Capital Investment Limit? $

Determine Susan's allowable QBI deduction. $
Answer #1

Calculation of QBI Limit

a. 20% of QBI i.e. (20%*2,00,000)=40000

or

b. 50% of the company’s W-2 wages OR the sum of 25% of the W-2 wages plus 2.5% of the unadjusted basis of all qualified property i.e (50% of 11000) = 5500 and (25% *11000)+ (2.5%*0) = 2750

Hence eligible deduction is Rs 2750

Net QBI = 200000-2750 =197250

Total Income= 197250+ 11000 (not to include interset income) = 208250

Hence, W-2 wages/ Capital Investment Limit = 11000/208250= 0.053